OpenStudy (anonymous):

solve −x^(2) − 5x < −6

OpenStudy (anonymous):

i have (x+6) (x-1) <0 is that correct so far?

OpenStudy (athe):

\[-x^2-5x+6<0\rightarrow x^2+5x-6>0\rightarrow (x-1)(x+6)>0\]

OpenStudy (eyust707):

yes that is very close.. you just forgot one little rule

OpenStudy (eyust707):

Whenever you multiply or divide both sides by a negative number you must flip the \(\lt\) sign around

OpenStudy (athe):

anwer \[\left( \infty, -6 \right)\cup\left( 1, \infty \right)\]

OpenStudy (anonymous):

eyust707 is athe correct?

OpenStudy (eyust707):

yes

OpenStudy (athe):

more correct \[\left( -\infty, -6 \right)\cup\left( 1, +\infty \right)\]

OpenStudy (eyust707):

ahh missed that

OpenStudy (eyust707):

I think you will be more familiar writng the solution like this: \[ x \gt 1 \space \space {\text or} \space \space x \lt -6 \]

OpenStudy (athe):

Yes, this is another record right answer.